A ball is projected vertically upward with a certain initial speed another ball of the same mass is projected at an angle of 60° with the vertical with the same initial speed at the highest point the ratio of their potential energy will be (I) 4:1 (ii) 3:2 (iii) 2:3 (iv) 2:1

Grade:11

2 Answers

Saswat Pani
6 years ago
In the horizontal motion E1= mgh1h1 =u×u/2gIn projectile motionE2=mgh2h2 = u×u×sin2theta÷2g = u×u×3/4÷2gRatio is H1/H2 which is 4/3
Aakash
6 years ago
In this question it is stated that the 2nd ball is thrown with angle 60° with the vertical that means 30° with the horizontal. Now just use: (u² sin²90/2g)/(u² sin²30/2g)= 1/1/4= 4/1Hence the ratio is 4:1Option (I)